### v2

# Both Ends Count! Just How Good are LLM Agents at Text-to-“Big SQL”?

Abstract.
Text-to-SQL and Big Data are both extensively benchmarked fields, yet there is limited research that evaluates them jointly. In the real world, Text-to-SQL systems are often embedded with Big Data workflows, such as large-scale data processing or interactive data analytics. We refer to this as “Text-to-Big SQL”. However, existing text-to-SQL benchmarks remain narrowly scoped and overlook the cost and performance implications that arise at scale. For instance, translation errors that are minor on small datasets lead to substantial cost and latency overheads as data scales, a relevant issue completely ignored by text-to-SQL metrics.
In this paper, we overcome this overlooked challenge by introducing novel and representative metrics for evaluating Text-to-Big SQL. Our study focuses on production-level LLM agents, a database-agnostic system adaptable to diverse user needs. Via an extensive evaluation of frontier models, we show that text-to-SQL metrics are insufficient for Big Data. In contrast, our proposed text-to-Big SQL metrics accurately reflect execution efficiency, cost, and the impact of data scale. Furthermore, we provide LLM-specific insights, including fine-grained, cross-model comparisons of latency and cost.
AI agents, Big Data, LLM, Text-to-SQL
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; April 27, 2026; Edinburgh, United Kingdomccs: Computing methodologies Natural language processingccs: Information systems Data analytics
1. Introduction
Text-to-SQL is a longstanding problem in NLP that seeks to bridge natural language (NL) interfaces and structured query generation. Recent advances in production Large Language Models (LLMs) have substantially improved cross-domain performance, placing them as effective text-to-SQL engines (anthropic2024claude3opus; openai2025gpt52; chung2025long-context) that can generalize across varying data schemas and terminology, improving state-of-the-art results (zhu2024largelanguagemodelenhanced; li2024dawn).
In practice, this generalization ability is further enhanced via the integration of AI agents, which act as task-specific LLM scaffolds to iteratively inspect database schemas, refine SQL generation, validate SQL syntax, enabling text-to-SQL execution to adapt to the unique traits of user-specific data sources (sapkota2026agents). Precisely, the existing ecosystem of open-source stacks for agent implementation (microsoft2024autogen; langchain2026; crewai2026home), combined with the current state of production LLMs (anthropic2024claude3opus; openai2025gpt52; chung2025long-context), has made text-to-SQL systems more accessible than ever.
However, when moving beyond traditional databases to Big Data systems, text-to-SQL faces additional complexities. For instance, systems such as Amazon Athena (aws_athena_overview) enable interactive analytics on massive datasets without requiring ETL, supporting various formats (CSV, JSON, ORC, Avro) and serverless, on-demand execution. In Big Data systems such as Athena, focusing only on the text-to-SQL end is not enough. The Big Data end itself introduces critical constraints that directly affect overall performance and cost. First, incorrect SQL has amplified consequences: failed queries can consume substantial compute resources, scan massive volumes of data, and increase execution costs, making accuracy essential not only for correctness but also for efficiency. For instance, (KoutsoukosM0K25) report that running TPC-H benchmark at a moderate scale factor of $100$ on Amazon Athena using Parquet data takes $132.3$ seconds, illustrating how failed queries can sharply increase execution time and costs in Big Data environments.
Second, inefficiencies can arise not only from failed queries but also from the SQL generation process itself. In agentic text-to-SQL, the LLM instructs the agent to use structured tools to inspect schemas and extract data-specific traits (deng2025reforcetexttosqlagentselfrefinement; xie2024magsqlmultiagentgenerativeapproach; zhang2024sqlfuseenhancingtexttosqlperformance). While this enables context-aware query adaptation, the reasoning overhead and tool orchestration can increase latency, making efficient SQL generation critical in Big Data systems. Simply put, if SQL generation becomes slower than physical query execution, interactive analysis may become impractical, undermining the performance gains of decades of optimizations in Query-as-a-Service (QaaS) engines like Athena and BigQuery (google_cloud_bigquery).
Overall, these reasons lead to the following observation:
{emphasizedBox}
Considering both SQL generation and execution is crucial, making text-to-SQL in Big Data environments distinct, for we refer to it as “Text-to-Big SQL. [^1] Here, “Big SQL” denotes a conceptual domain in SQL query processing; it is unrelated to IBM’s Db2 Big SQL product (ibm_db2_big_sql). “Big SQL” is a trademark of IBM Corporation, used descriptively and with no affiliation.
1.1. Why Current Text-to-SQL Approaches Fall Short for Big Data?
Traditional text-to-SQL methods have largely been evaluated on moderate-scale relational databases, focusing on query translation or isolated accuracy metrics (spider202025; li2024dawn; zhang2024finsql; deochake2025costawaretexttosqlempiricalstudy). While these benchmarks provide insights into LLM capabilities, they often overlook the complexities of interactive execution, streaming data, and cost considerations (cheng2025barbariansgateaiupending).
One example of this is that many text-to-SQL benchmarks use binary correctness metrics, tagging each generated query with a simple 0/11 label, thereby diluting degrees of partial correctness (li2023bird; zhang2024finsql). This is obvious, for example, when a generated query incorrectly projects an unnecessary column. In traditional text-to-SQL this counts as a wrong translation, but in text-to-Big SQL it should be partially acceptable, since re-running a query for a single extra column could be extremely costly. Consequently, new evaluation metrics are needed to jointly account for partial correctness and cost, reflecting practical text-to-SQL performance for Big Data.
While LLMs are effective text-to-SQL processors (li2025OmniSQL; li2024dawn), their performance depends on how agents scaffold tool use (sapkota2026agents; yao2023react). In a ReAct-style framework, the LLM controller guides reasoning, selects tools, and interprets feedback, while the executor runs the tools. Fast tools, such as fetching a table schema from a data catalog can be bottlenecked by extensive LLM reasoning for query validation, or vice versa. Efficient interaction between LLM, agent, and tools is thus critical for responsive interactive analytics. However, there are no evaluations that focus on this interplay and its effect on Big Data performance. In this paper, we start addressing this gap by jointly measuring agent and tool utilization, reasoning latency, and downstream query execution.
1.2. Our Contribution
In this work, we propose a benchmarking methodology for text-to-Big SQL agents that treats both ends, namely query generation and execution, as first-class citizens. We focus on zero-shot LLM agents to examine a worst-case scenario where no specific fine-tuning or additional optimization is applied, revealing the true impact of SQL generation, agent action, and tool interaction on performance and cost in Big Data settings. Our contributions are the following:
  1. (1)
    A novel evaluation framework for text-to-SQL agents designed to capture big query execution. We propose new metrics that jointly assess agent action, reasoning latency, and the cost-effectiveness of generated queries, reflecting both partial correctness and the practical implications of running queries on Big Data engines.
  2. (2)
    A systematic evalution of state-of-the-art LLMs within a unified ReAct-style agent architecture. This analysis reveals insights beyond accuracy, identifying scenarios where newer models achieve high correctness but are less interactive due to reasoning or tool orchestration overhead, which happens for instance with Opus 4.6.
  3. (3)
    A discussion of the unique challenges, as well as open research questions in text-to-Big SQL, including the interplay between SQL generation, agent tool use, and execution performance, an area largely overlooked by existing benchmarks.
2. Text-to-Big SQL Demands New Metrics
2.1. Limitations of Current Metrics
Typically, a text-to-SQL benchmark suite comprises a set of triples containing a natural language (NL) query, a golden query in SQL, and a ground truth ($V^{n}$) result (li2023bird). During evaluation, system performance is measured by comparing the generated SQL to the golden query, and the resulting output ($\hat{V}^{n}$) to $V^{n}$. Overall accuracy is then computed by aggregating these comparisons across the benchmark suite using standard evaluation metrics, such as Exact Matching (EM), Execution Accuracy (EA), and Valid Efficiency Score (VES) (hong2025database-interfaces; zhu2024largelanguagemodelenhanced; luo2025state-of-the-art).
The major limitation of these metrics is their reliance on all-or-nothing correctness [^2] The corresponding formulas are provided in the Appendix. And extended version of this work is available on arXiv: https://arxiv.org/abs/2602.21480. In practice, however, a generated query may deviate from the expected result without being entirely invalid, provided that end users may still be able to determine its correctness through simple validation, for example, by detecting an additional column in the projected output. We summarize the possible outcomes of an SQL execution as follows:
A
Incorrect row count: The result is invalid due to poor SQL translation, for example, mis-specified WHERE conditions or inappropriate join types (INNER, LEFT, etc.), leading to an incorrect number of rows. Such silent failures may remain unnoticed by the user, but always need a new translation and re-execution cycle to ensure correctness.
B
Missing columns: The result is invalid due to omission of mandatory attributes in the projection list, requiring query re-execution to retrieve the complete output by adding the missing attributes in the SELECT clause.
C
Superfluous columns: The result is valid, as we assume that experienced users can manually drop the extra columns without modifying the returned output, which is very cheap and fast. For instance, df.drop(‘‘extra_col’’) returns a new DataFrame without the extra column in Spark (ApacheSparkSQL). However, processing superfluous data affects execution performance and cost and should be penalized in the assessment.
2.2. Proposed Metrics
To encode our notion of query correctness in a measurable metric, we extend the standard text-to-SQL VES metric to account for superfluous columns. Specifically, to quantify the overhead introduced by including irrelevant columns, we compute the column-level precision: $P(S,\hat{S})=\frac{S\cap\hat{S}|}{|\hat{S}|}$, where $S$ is the set of ground-truth columns and $\hat{S}$ is the set of columns in the execution result. This captures the fraction of retrieved columns that are actually relevant, penalizing additional, unnecessary columns without discarding partially correct results ( C).
Also, our new metric considers the total end-to-end (e2e) time ($T_{\text{e2e}}$), which includes all back-and-forth interactions between the LLM and the agent, the execution of interim tools, as well as the time required to run the generated SQL query on the underlying Big Data engine. Combining both concepts, we propose the novel text-to-Big SQL metric called $\mathrm{VES}^{*}$, which is defined as follows for $N$ queries:
(1) $\mathrm{VES}^{*}=\frac{1}{N}\sum_{i=1}^{N}\left(\mathds{1}(V_{i},\hat{V}_{i})\cdot P(S_{i},\hat{S}_{i})\cdot\frac{T_{gold}}{T_{e2e}}\right),$
where $T_{gold}$ denotes the execution time of the golden query, and $\mathds{1}(V_{i},\hat{V}_{i})$ is an indicator function that tells whether the output of the generated query $\hat{V}_{i}$ matches the expected result $V_{i}$ ( A and B):
(2) $\mathds{1}(V,\hat{V})=\begin{cases}1,&\text{if $V$ is contained in expected output $\hat{V}$}\\ 0,&\text{otherwise}\end{cases},$
We also introduce the Valid Cost-Efficiency Score (VCES), a cost-oriented derivative of $\text{VES}^{*}$ to account for the overall execution cost $C_{\text{e2e}}$, including the iterative interactions between the LLM and agent, the execution of agent-invoked tools, and the runtime of the generated query on the Big Data engine. For text-to-Big SQL, benchmarking query execution cost is relevant, particularly in cloud deployments:
(3) $\text{VCES}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{\mathbf{1}(V_{i},\hat{V}_{i})\cdot P(S_{i},\hat{S}_{i})\cdot\frac{T_{gold}}{T_{e2e}}}{C_{e2e}}\right).$
To complement the previous metrics, we introduce the Expected Cost per Valid Query (CVQ), which quantifies the anticipated cost to obtain a valid result under a retry-until-success strategy. Let $p$ denote the single-shot validity rate, that is, the fraction of generated big queries that are valid. Consequently, the expected number of attempts until success follows a geometric distribution with mean $1/p$. Accordingly, we define $CVQ=\frac{C_{\text{e2e}}}{p}$.
3. AI Agent Design
To evaluate the novel text-to-Big SQL paradigm, we use a ReAct (Reasoning + Acting) agent (yao2023react), a well-established framework (sapkota2026agents; paperbench2025; liu2025palimpchat), which decomposes agentic operation into three intertwined components: Thought, Action, and Observation. Via Thought, the agent reasons about the task and decides its next step; through Action, it interacts with the environment or an external tool to gather or process information; and through Observation, it interprets feedback from these actions to refine subsequent reasoning steps.
We deliberately keep the agent simple to provide a broader perspective in our analysis. While more complex AI agents exist (li2025alphasqlzeroshottexttosqlusing), we opt to leverage the long-context capabilities of production LLMs guided by iterative cycles of Thought, Action, and Observation, which has already demonstrated its effectiveness in text-to-SQL tasks (chung2025long-context).
We define the controller as the connected LLM that guides ReAct-style reasoning, decides when and how to use tools, and produces the final answer. Conversely, the executor is the program that connects the LLM with external tools and handles the execution loop. We illustrate the architecture and execution loop of the agent in the Appendix.
For the downstream query engine, we choose Spark SQL (ApacheSparkSQL) due to its widespread adoption for large-scale structured data analysis. The controller interacts with the Spark session through a set of four tools, whose specifications are fully inserted into the context in a zero-shot manner (zero-shot-tools2023). We define the tools as follows:
1:
list_tables: Look up the Spark Catalog (pyspark_catalog_docs) to retrieve available tables by executing a SHOW TABLES statement.
2:
get_schema: Retrieve the schema for one or more tables from the Spark Catalog using SHOW CREATE TABLE t. Optionally, the controller can fetch an adjustable number of sample rows from each table with the same tool call, which triggers an additional SELECT * FROM t statement.
3:
check_query: Verify the syntax of a proposed query using predefined heuristics (chung2025long-context) using a connected LLM (the checker). While we use the same LLM for both the checker and the controller, they can be different.
4:
run_query: Execute a SQL query in the connected Spark Session, regardless of whether the deployment is local or cluster-based.
To prevent the model from iteratively correcting and re-running queries, we terminate the agent immediately after the first run_query execution. We take this design decision because, in Big Data systems, unrestricted execution loops can lead to excessive resource consumption or high billing costs, and may even result in stuck-in-the-loop (cheng2025barbariansgateaiupending) scenarios without improving the accuracy of the inferred query.
We base our implementation on the original LangChain Spark SQL Toolkit (langchain_sparksql_2026) but port it to LangGraph (langchain_langgraph_2024). We utilize the LangChain stack because it is actively maintained (langchain2026) and frequently serves as a research baseline (ding2025unifiedtoolintegrationllms; bhagat-etal-2025-evaluating; heng2025langchain-hpc).
4. Evaluation
Our evaluation proceeds as follows. §4.1 shows that text-to-SQL metrics lack informativeness, revealing opportunities from fine-grained agent benchmarking. §4.2 demonstrates the superior discriminability of text-to-Big SQL versus text-to-SQL metrics in differentiating agents across latency and cost objectives. Finally, §4.3 shows data scale matters and that its impact can be quantified via text-to-Big SQL metrics. We employ two benchmarks: a text-to-SQL-focused one and a Big Data-focused one.
  • BIRD (li2023bird), a text-to-SQL benchmark that assesses translation accuracy for realistic databases.
  • TPC-H (tpc_h_spec), a classic data analytics benchmark that measures database performance on complex ad-hoc business queries over relational data. TPC-H is very useful for text-to-Big SQL because it allows for the deterministic scaling of data.
We conducted all experiments on an AWS m5.xlarge EC2 (aws_ec2_product) instance in the us-east-1 region. We selected a set of representative frontier models from various providers for the evaluation. For all Large Language Models (LLMs), we used the official provider APIs and calculated costs based on their respective per-token pricing.
To ensure consistency for interactive use, all models are deployed with low-latency reasoning configurations. We standardize sampling hyperparameters, such as temperature, top-$p$, and maximum token limits, across all APIs, except where specific parameters are not exposed by the provider. These configurations are detailed in the Appendix.
4.1. Accuracy is Not Enough in Text-to-Big SQL
When models achieve similar accuracy, standard text-to-SQL metrics fail to differentiate between setups. We prove this by inspecting the Execution-based Focused Evaluation (EX) (spider202025), a SOTA text-to-SQL accuracy metric. Since text-to-SQL performance continues to improve independently (hong2025database-interfaces), we design a practical testbed assuming near-perfect accuracy. To this end, we select eight queries from the BIRD dataset where all tested models (except the later-added GPT-5.2) attain an average EX of at least 0.85.
EX and e2e execution time alone lack the informativeness to discriminate models effectively. Figure 1 shows this: for faster models, the accuracy/speed tradeoff is unclear: e.g., is GPT-4o better than Gemini 3 Flash (27.79% faster but imperfect accuracy)? If wrong-query costs are high, a slightly slower but more accurate model may be preferable in production.
Later-generation models do not clearly outperform their predecessors in zero-shot agentic text-to-SQL. For example, Opus 4.6 achieves perfect accuracy but takes 92.37% longer execution time than GPT-4o. Similarly, Gemini 3 Pro exhibits poor latency, further exacerbated by high variance from API instabilities. Notably, GPT-4o was released in 2024, nearly two years before the other two models.
Table 1. EX, total execution time and the breakdown of execution time per ”Observation-Thought-Action” stage for each LLM, on selected BIRD queries. Results are displayed in descending order of total execution time.
Model EX E2E (s) % of E2E Time
Mean $\sigma$ list schema check run
GPT-4o 0.93 6.55 2.08 9.22 13.15 62.08 13.64
Gemini 3 Flash 1.00 8.37 2.61 9.74 10.15 66.76 11.71
GPT-5.2 0.69 8.44 2.74 13.18 21.36 48.88 14.88
Gemini 2.5 Flash 0.95 9.18 3.21 11.70 10.53 66.07 10.11
Claude Opus 4.5 1.00 11.40 2.82 16.90 18.57 42.32 20.93
Claude Opus 4.6 1.00 12.60 2.09 17.18 18.11 42.70 20.93
Kimi K2.5 0.98 13.61 5.72 9.99 14.35 62.89 11.74
GPT-5 0.88 15.45 8.63 7.91 11.09 72.04 7.98
GLM-5 1.00 79.63 50.57 11.77 10.60 66.91 10.53
Gemini 3 Pro 1.00 115.95 159.97 18.29 19.70 46.82 14.33
For better observability, we breakdown execution time, aggregating Observation-Thought-Action iterations that call the same tool into the stage abstraction. Common patterns emerge across models: the check_query stage dominates end-to-end time in all LLMs, as expected since it runs within the LLM rather than the local Spark session. Yet percentages vary widely (e.g., a 23% spread between GPT-5 and GPT-5.2).
These results suggest stage-specific optimization via model selection. For instance, the time split between list_tables and run_query differs by model. Overall, smart per-stage model assignment (as in model ensembles (cheng2025barbariansgateaiupending)) offers clear optimization potential.
4.2. Big SQL Metrics Zoom In on Performance
In the context of Big Data, text-to-SQL metrics fail to inform model selection. Instead, a Big SQL lens differentiates LLMs more clearly, providing sharper selection criteria. Table 2 shows normalized VES and VES* with respect to the best scoring LLM setup: GPT-4o. As shown in the table, the VES* metric better discriminates accurate models (809.09% dispersion range versus 54.93% in VES). The main reason is that VES considers only query execution time and result accuracy (hong2025database-interfaces). However, modern LLMs excel at text-to-SQL: their generated queries often functionally resemble the gold query (even if not identical) yielding similar execution times. At high accuracy levels, this hinders model discrimination. Our proposed metric addresses this limitation by breaking ties, favoring LLMs that combine efficient agentic interaction with minimal projection overhead (i.e., avoiding superfluous columns).
Table 2. VES and VES* for the selected BIRD queries. Both metrics are normalized to the best scoring LLM. LLMs are displayed in descending order of VES*.
Model VES (norm) VES* (norm) Time Variation (x)
list schema check run
GPT-4o 1.00 1.00 1.00x 1.00x 1.00x 1.00x
Gemini 3 Flash 1.06 0.81 1.35x 0.99x 1.37x 1.10x
Gemini 2.5 Flash 1.00 0.78 1.78x 1.12x 1.49x 1.04x
Claude Opus 4.5 1.09 0.57 3.19x 2.46x 1.19x 2.67x
Claude Opus 4.6 1.09 0.51 3.58x 2.65x 1.32x 2.95x
GPT-5 0.89 0.46 2.02x 1.99x 2.74x 1.38x
Kimi K2.5 1.05 0.45 2.25x 2.27x 2.10x 1.79x
GPT-5.2 0.71 0.39 1.84x 2.09x 1.01x 1.41x
Gemini 3 Pro 1.10 0.33 35.12x 26.52x 13.35x 18.61x
GLM-5 1.05 0.11 15.52x 9.81x 13.10x 9.39x
A high VES* effectively reflects both accurate and low latency models, ranking GPT-4o highest (Table 2). It also reveals nuances, such as the Opus models benefiting from perfect accuracy despite incurring higher execution times. When normalizing execution times relative to the fastest model (GPT-4o), GPT-4o consistently “wins” across all stages, indicating a clear separation between“fast” and “slow” model classes. VES is unable to capture this system behavior.
Table 3. VCES and CVQ for the selected BIRD queries. The VCES values are normalized to the best-performing LLM, and models are shown in descending order of VCES.
Model VCES norm. ($\mathdollar^{-1}$) CVQ (\$) Cost Variation (x)
list schema check run
Gemini 3 Flash 1.00 0.0044 1.00x 1.00x 1.00x 1.00x
Gemini 2.5 Flash 0.85 0.0053 1.36x 1.18x 1.12x 0.98x
GPT-4o 0.55 0.0107 2.14x 2.93x 2.10x 2.64x
Kimi K2.5 0.53 0.0047 1.07x 1.48x 0.99x 1.05x
GPT-5.2 0.25 0.0124 2.62x 4.07x 1.42x 2.46x
GPT-5 0.24 0.0118 1.92x 2.58x 2.55x 1.61x
Claude Opus 4.5 0.08 0.0388 15.30x 16.13x 5.59x 15.76x
Claude Opus 4.6 0.08 0.0359 14.39x 14.56x 5.22x 14.58x
Gemini 3 Pro 0.07 0.0254 10.85x 11.21x 4.05x 7.07x
GLM-5 0.05 0.0129 3.54x 3.06x 2.94x 2.63x
VCES matches the discriminative power of VES* while incorporating cost, as it factors in per-token billing and the expenses of suboptimal query execution. Table 3 presents VCES per LLM and identifies Gemini 3 Flash as the most cost-efficient option due to its low per-token pricing3330.5/3.0/3.0 per input/output token for Gemini 3 Flash (GoogleGeminiPricing2026), vs. 2.5/10.0/10.0 for GPT-4o (OpenAIPricing2026).. Ultimately, VCES complements VES* by enabling the selection of cost-efficient LLM setups, thus addressing a significant gap in text-to-SQL metrics that directly ignore cost. We attach CVQ to better support our finding: GPT-4o more than doubles the per-query cost of Gemini 3 Flash due to its lower accuracy, which leads to more failed queries and higher token consumption.
As with execution time results, the leading cost performer in Table 3 dominates across nearly all stages. Interestingly, the results reveal a clear trade-off between latency-optimal and cost-efficient models: for example, GPT-4o ranks first in execution speed but achieves only half the cost-efficiency of Gemini 3 Flash, which is slightly slower but much cheaper. This distinction suggests that cost-efficient models may be preferable for stages with relaxed latency requirements, while VES*-optimal models are ideal when latency is critical.
4.3. The Aftermath of Data Scale
[…]
Figure 1. Breakdown of agent execution time (a) and cost (b) across different scale factors for TPC-H Query 21. Each bar series represents a specific model: Gemini 3 Pro (G), Claude Opus 4.5 (A), and GPT-5.2 (O).
The remaining key question is the extent to which a new metric is needed to capture data scaling effects. To examine this, we chose four TPC-H queries where LLMs deliver poor accuracy (Figure 2): three complex ones (Q17, Q18, Q21) with nested subqueries, plus one simple single-table query (Q1). We used TPC-H business questions as NL inputs and official SQL as ground truth (see Appendix for details). Tests ran on Amazon EMR (aws_emr_2026) with r5b.xlarge master/core nodes, 32 core nodes (4 vCPUs each), and 128 GB gp3 EBS volumes.
Figure 2. Execution accuracy across three models for TPC-H queries 1, 17, 18, and 21. Accuracy is computed as the average value of Equation 2 over 10 zero-shot translations of each corresponding natural language query.
As illustrated in Figure 3. VES (which incorporates SQL execution time) yields constant relationships between models across scale factors. In contrast, CVQ better captures each model’s potential cost loss at varying scales: less accurate models (e.g., Gemini 3 Pro here) pose greater risk at large scale factors, where query errors incur substantially higher costs. Consequently, failing a query at SF 1000 is far more expensive than at SF 10. Even a modest accuracy gap (such as the 10% difference between Opus 4.5 and GPT-5.2) becomes critically amplified at higher scales.
(a) VES
(b) CVQ
Figure 3. Text-to-SQL (a) and text-to-Big SQL (b) metrics across three models and scale factors, averaged for TPC-H queries 1, 17, 18, and 21.
Our evaluation clarifies the interpretation of text-to-Big SQL metrics. VES* and VCES provide practical assessments at fixed data scale factors, as they normalize generated SQL execution time and cost to the ground truth. In contrast, CVQ serves as an essential complement by quantifying the amplified impact of query inaccuracies as data scales.
5. Discussion and Future Opportunities
Based on our result, we identify several promising avenues in the broader text-to-Big SQL domain. Figure 4 summarizes our key insight: even if text-to-SQL reached 99% accuracy since natural language ambiguity is unavoidable, text-to-Big SQL challenges would still remain.
Figure 4. Challenges of Text-to-Big SQL include, but are not limited to, those of Text-to-SQL.
Agent performance tuning. Optimizing the internal stages of agents itself presents a significant research opportunity. Both our findings and prior work (cheng2025barbariansgateaiupending) demonstrate that both performance and cost could be improved by “strategically” assigning specialized models to different stages (e.g., navigating “fast” and‘ “cheap” models). However, existing physical plan optimization approaches for semantic operators incur latencies of tens of seconds (Russo_Liu_Sudhir_Vitagliano_Cafarella_Kraska_Madden_2026; zhu2025relationalsemanticawaremultimodalanalytics), making them incompatible with interactive analytics. Adapting these models to meet the latency constraints of interactive Big Data analytics remains a major open challenge.
Text-to-SQL must be optimized for large-scale query execution. In Big Data, syntactically correct SQL may still be impractical if it triggers large shuffles, unnecessary joins, or full-table scans. Text-to-Big SQL must therefore optimize for both correctness and cost-efficient, large-scale execution, which remains an open challenge.
One promising direction is to leverage historical execution traces enriched with performance metrics and system-level quality indicators, such as VCES and CVQ. By semantically matching newly generated queries (fu2023catsql) to past executions (wang2025andromeda), the system can estimate expected cost and runtime before execution and proactively rewrite inefficient queries (song2026quitequeryrewriterules; he2025aqorafastlearnedadaptive). Incorporating physical plans and cost models (baldacci2019cost_model) further enables extrapolation across data scales, supporting scale-aware optimization rather than static SQL translation. Alternatively, the agent can be few-shot with similar past queries and their Big SQL metrics to infer optimized SQL.
Beyond exact execution, Big Data systems often rely on approximate queries to trade precision for performance (chaudhuri17). However, current text-to-SQL models rarely reason about such semantic alternatives. A text-to-Big SQL framework should instead consider approximate joins, sampling-based aggregations, or sketch-based summaries when they satisfy user intent while significantly reducing execution cost. User-provided QoS annotations (e.g., “run this fast”) could further guide optimization.
User-defined functions (UDFs). Another key challenge in text-to-Big SQL arises from UDFs. Big Data engines like Spark, Athena, and BigQuery often leverage custom UDFs, which are not fully expressible in standard SQL. As a result, text-to- Big SQL solutions must produce UDF-compatible SQL or hybrid code, for example, combining SQL with Spark DataFrames, which goes beyond classical text-to-SQL.
6. Conclusion
In this work, we have started addressing the surprisingly underexplored domain of “text-to-Big SQL”: the integration of text-to-SQL systems within Big Data engines. We have shown that standard text-to-SQL benchmarks fall short and, for the first time, introduced text-to-Big SQL measures that reflect the true interaction between agents, LLMs, Big Data systems, and data scale. We have evaluated state-of-the-art production LLMs to demonstrate that our metrics provide an effective assessment framework for text-to-Big SQL. Our work exposes new real-world challenges and guides future research in text-to-Big SQL.
Acknowledgements.
This work has been partly funded by the EU Horizon programme under the grant agreements: 101092644 (NearData), 101092646 (CloudSkin), 101093110 (EXTRACT) and 101086248 (CLOUDSTARS). Germán T. Eizaguirre is recipient of a pre-doctoral FPU grant from the Spanish Ministry of Universities (ref. FPU21/00630). Marc Sánchez-Artigas is a Serra Húnter Fellow.
Appendix
Appendix A Agent Architecture
Figure 5. The architecture of our agent is centered on an executor built with LangGraph. For our evaluation, both the controller and the checker operate on the same LLM. The Spark Session can be connected to either a local deployment or a distributed cluster.
Appendix B Related Work
While other works address Big Data in text-to-SQL, they often overlook downstream analysis. Existing benchmarks focus on complex upstream data sources (spider202025; li2024dawn; zhang2024finsql) or isolate execution cost from accuracy (deochake2025costawaretexttosqlempiricalstudy; zhang2024finsql). Others study LLM performance at scale but emphasize data diversity over job cost (li2025OmniSQL). We target the gap in this domain, which is already a reality in production with tools like BigQuery’s Generative AI (google2026bigquerygenai).
Non-binary accuracy metrics exist (spider202025; Pinna2025redefining), and some works apply similar logic to improve column linking (knapsack2025). However, we are the first to integrate this idea into the Big Data domain to account for their influence on SQL execution performance.
LLMs are effective text-to-SQL processors (li2025OmniSQL; li2024dawn), yet they require agent-based scaffolding to interact with user-specific databases (luo2025state-of-the-art). While some agents integrate with Big Data (heng2025langchain-hpc), their text-to-SQL effectiveness remains unevaluated. Conversely, current agent benchmarks (huo2025birdinteractreimaginingtexttosqlevaluation) neglect the query execution. This work introduces an evaluation framework that jointly assesses agent interactivity, accuracy and job performance.
Appendix C Text-to-SQL formulas
A text-to-SQL benchmark suite includes a set of triples containing a natural language (NL) query, a golden query in SQL ($Q^{n}$), and a ground truth ($V^{n}$) result (li2023bird). During evaluation, system performance is measured by comparing the generated SQL ($\hat{Q}^{n}$) to the golden query, and the resulting output ($\hat{V}^{n}$) to $V^{n}$ (hong2025database-interfaces).
C.1. Exact Matching (EM)
(4) $EM=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(Q_{i},\hat{Q}_{i}\right)$
where
(5) $\mathbb{I}(Q_{i},\hat{Q}_{i})=\begin{cases}1,&Q_{i}=\hat{Q}_{i}\\ 0,&Q_{i}\neq\hat{Q}_{i}\end{cases}$
C.2. Execution Accuracy (EA)
(6) $EA=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(V_{i},\hat{V}_{i}\right)$
where
(7) $\mathbb{I}(V_{i},\hat{V}_{i})=\begin{cases}1,&V_{i}=\hat{V}_{i}\\ 0,&V_{i}\neq\hat{V}_{i}\end{cases}$
C.3. Valid Efficiency Score
(8) $\text{VES}=\frac{1}{N}\sum_{i=1}^{N}\left(\mathbb{I}(V_{i},\hat{V}_{i})\cdot\frac{T_{gold}}{T_{gen}}\right)$
Where $T_{gold}$ and $T_{gen}$ are the execution times of the golden query and the generated SQL, respectively.
Appendix D BIRD evaluation
D.1. Model Selection
We selected the models according to the following criteria: (1) current frontier models from Google, Anthropic, and OpenAI; (2) previous-generation frontier models from these same providers; and (3) the three open-source models with the highest scores on SWE-rebench  [^4] https://swe-rebench.com/ (all at the time of the experiment, mid-February 2026).
D.2. Detailed Results
Our BIRD evaluation averages each metric over 50 iterations. Table LABEL:tab:appendix-bird-queries provides a detailed breakdown of the resulting Text-to-SQL and Text-to-Big SQL metrics per query and model.
Table 4. Per-query BIRD metrics by model (mean $\pm$ std across runs). VES, VES*, and VCES follow the paper definitions; SVQ reports the expected cost per valid query.
Query ID Model EX VES VES* VCES SVQ
61 GPT-4o 1.00 $\pm$ 0.00 1.0776 $\pm$ 0.1609 0.0111 $\pm$ 0.0040 1.1146 $\pm$ 0.4623 0.0099 $\pm$ 0.0008
GPT-5.2 1.00 $\pm$ 0.00 0.9884 $\pm$ 0.1617 0.0063 $\pm$ 0.0022 0.6872 $\pm$ 0.3280 0.0092 $\pm$ 0.0019
Claude Opus 4.6 1.00 $\pm$ 0.00 1.0577 $\pm$ 0.1326 0.0060 $\pm$ 0.0017 0.2045 $\pm$ 0.0562 0.0296 $\pm$ 0.0002
Gemini 3 Flash 1.00 $\pm$ 0.00 1.0612 $\pm$ 0.1561 0.0056 $\pm$ 0.0019 0.9073 $\pm$ 0.5326 0.0062 $\pm$ 0.0016
Claude Opus 4.5 1.00 $\pm$ 0.00 1.0707 $\pm$ 0.1468 0.0055 $\pm$ 0.0016 0.1368 $\pm$ 0.0445 0.0401 $\pm$ 0.0033
Kimi K2.5 1.00 $\pm$ 0.00 1.0837 $\pm$ 0.1747 0.0052 $\pm$ 0.0013 1.1084 $\pm$ 0.3150 0.0046 $\pm$ 0.0004
DeepSeek Chat 0.86 $\pm$ 0.35 0.9004 $\pm$ 0.4139 0.0033 $\pm$ 0.0018 1.7322 $\pm$ 1.4172 0.0022 $\pm$ 0.0002
Gemini 3 Pro 1.00 $\pm$ 0.00 1.0892 $\pm$ 0.1364 0.0031 $\pm$ 0.0011 0.1487 $\pm$ 0.0728 0.0210 $\pm$ 0.0040
Gemini 2.5 Flash 0.60 $\pm$ 0.49 0.6223 $\pm$ 0.5169 0.0031 $\pm$ 0.0030 0.4032 $\pm$ 0.5152 0.0128 $\pm$ 0.0017
GPT-5 1.00 $\pm$ 0.00 1.0561 $\pm$ 0.1747 0.0029 $\pm$ 0.0009 0.2275 $\pm$ 0.0848 0.0129 $\pm$ 0.0017
GLM-5 1.00 $\pm$ 0.00 1.0153 $\pm$ 0.1511 0.0008 $\pm$ 0.0004 0.0531 $\pm$ 0.0435 0.0150 $\pm$ 0.0056
606 Gemini 3 Flash 1.00 $\pm$ 0.00 0.9771 $\pm$ 0.0830 0.0115 $\pm$ 0.0021 1.8689 $\pm$ 0.6777 0.0062 $\pm$ 0.0011
Claude Opus 4.5 1.00 $\pm$ 0.00 1.0093 $\pm$ 0.0652 0.0114 $\pm$ 0.0012 0.3002 $\pm$ 0.0452 0.0379 $\pm$ 0.0012
Gemini 2.5 Flash 1.00 $\pm$ 0.00 0.9724 $\pm$ 0.0780 0.0112 $\pm$ 0.0032 1.6157 $\pm$ 1.1519 0.0069 $\pm$ 0.0027
GPT-4o 0.52 $\pm$ 0.50 0.5069 $\pm$ 0.4884 0.0106 $\pm$ 0.0114 1.0953 $\pm$ 1.3342 0.0187 $\pm$ 0.0013
Claude Opus 4.6 1.00 $\pm$ 0.00 0.9840 $\pm$ 0.0805 0.0103 $\pm$ 0.0009 0.2953 $\pm$ 0.0259 0.0349 $\pm$ 0.0002
Kimi K2.5 1.00 $\pm$ 0.00 0.9998 $\pm$ 0.0674 0.0097 $\pm$ 0.0016 2.1353 $\pm$ 0.4786 0.0046 $\pm$ 0.0005
DeepSeek Chat 0.92 $\pm$ 0.27 0.8326 $\pm$ 0.3136 0.0060 $\pm$ 0.0019 2.8653 $\pm$ 0.9997 0.0023 $\pm$ 0.0002
GLM-5 1.00 $\pm$ 0.00 0.9554 $\pm$ 0.0817 0.0012 $\pm$ 0.0006 0.0580 $\pm$ 0.0723 0.0215 $\pm$ 0.0087
Gemini 3 Pro 1.00 $\pm$ 0.00 0.9385 $\pm$ 0.1275 0.0012 $\pm$ 0.0018 0.0358 $\pm$ 0.1153 0.0342 $\pm$ 0.0108
GPT-5.2 0.04 $\pm$ 0.20 0.0212 $\pm$ 0.1049 0.0007 $\pm$ 0.0033 0.0849 $\pm$ 0.4161 0.1979 $\pm$ 0.0002
GPT-5 0.06 $\pm$ 0.24 0.0312 $\pm$ 0.1238 0.0003 $\pm$ 0.0014 0.0228 $\pm$ 0.1033 0.2526 $\pm$ 0.0023
645 GPT-4o 1.00 $\pm$ 0.00 0.9775 $\pm$ 0.1205 0.0115 $\pm$ 0.0030 1.3142 $\pm$ 0.4871 0.0088 $\pm$ 0.0011
GPT-5.2 1.00 $\pm$ 0.00 0.9913 $\pm$ 0.1306 0.0111 $\pm$ 0.0017 1.9508 $\pm$ 0.3330 0.0057 $\pm$ 0.0002
Gemini 3 Flash 1.00 $\pm$ 0.00 0.9947 $\pm$ 0.0889 0.0097 $\pm$ 0.0012 2.8552 $\pm$ 0.5324 0.0034 $\pm$ 0.0003
Gemini 2.5 Flash 1.00 $\pm$ 0.00 0.9907 $\pm$ 0.1083 0.0091 $\pm$ 0.0013 2.3973 $\pm$ 0.5125 0.0038 $\pm$ 0.0004
Kimi K2.5 1.00 $\pm$ 0.00 0.9993 $\pm$ 0.0927 0.0060 $\pm$ 0.0007 1.5733 $\pm$ 0.2495 0.0038 $\pm$ 0.0002
Claude Opus 4.5 1.00 $\pm$ 0.00 0.9909 $\pm$ 0.0996 0.0058 $\pm$ 0.0008 0.1549 $\pm$ 0.0223 0.0374 $\pm$ 0.0007
GPT-5 1.00 $\pm$ 0.00 0.9931 $\pm$ 0.1166 0.0057 $\pm$ 0.0013 0.7266 $\pm$ 0.1910 0.0078 $\pm$ 0.0004
Claude Opus 4.6 1.00 $\pm$ 0.00 1.0188 $\pm$ 0.1081 0.0055 $\pm$ 0.0007 0.1604 $\pm$ 0.0214 0.0341 $\pm$ 0.0002
DeepSeek Chat 1.00 $\pm$ 0.00 0.9986 $\pm$ 0.0854 0.0037 $\pm$ 0.0003 1.9634 $\pm$ 0.2096 0.0019 $\pm$ 0.0001
GLM-5 1.00 $\pm$ 0.00 0.9793 $\pm$ 0.0789 0.0014 $\pm$ 0.0004 0.1458 $\pm$ 0.0576 0.0094 $\pm$ 0.0013
Gemini 3 Pro 1.00 $\pm$ 0.00 0.9824 $\pm$ 0.1100 0.0004 $\pm$ 0.0003 0.0127 $\pm$ 0.0690 0.0323 $\pm$ 0.0128
776 GPT-4o 0.98 $\pm$ 0.14 0.9596 $\pm$ 0.2170 0.0090 $\pm$ 0.0032 0.8559 $\pm$ 0.3818 0.0107 $\pm$ 0.0013
Gemini 3 Flash 1.00 $\pm$ 0.00 0.9893 $\pm$ 0.1598 0.0087 $\pm$ 0.0017 2.1869 $\pm$ 0.5540 0.0040 $\pm$ 0.0006
Gemini 2.5 Flash 1.00 $\pm$ 0.00 0.9971 $\pm$ 0.1216 0.0083 $\pm$ 0.0014 1.9412 $\pm$ 0.4800 0.0043 $\pm$ 0.0005
GPT-5.2 1.00 $\pm$ 0.00 1.0895 $\pm$ 0.1598 0.0064 $\pm$ 0.0013 0.6279 $\pm$ 0.2007 0.0102 $\pm$ 0.0012
Claude Opus 4.5 1.00 $\pm$ 0.00 1.0796 $\pm$ 0.1737 0.0058 $\pm$ 0.0009 0.1463 $\pm$ 0.0236 0.0398 $\pm$ 0.0002
Claude Opus 4.6 1.00 $\pm$ 0.00 1.1299 $\pm$ 0.1927 0.0051 $\pm$ 0.0011 0.1362 $\pm$ 0.0292 0.0373 $\pm$ 0.0001
Kimi K2.5 0.96 $\pm$ 0.20 0.9410 $\pm$ 0.2426 0.0046 $\pm$ 0.0014 0.9439 $\pm$ 0.3565 0.0050 $\pm$ 0.0008
GPT-5 1.00 $\pm$ 0.00 1.0371 $\pm$ 0.2007 0.0044 $\pm$ 0.0010 0.4871 $\pm$ 0.1361 0.0091 $\pm$ 0.0008
DeepSeek Chat 0.50 $\pm$ 0.50 0.4874 $\pm$ 0.5029 0.0016 $\pm$ 0.0017 0.7707 $\pm$ 0.7466 0.0042 $\pm$ 0.0001
GLM-5 1.00 $\pm$ 0.00 1.1182 $\pm$ 0.1901 0.0013 $\pm$ 0.0003 0.1146 $\pm$ 0.0379 0.0112 $\pm$ 0.0027
Gemini 3 Pro 1.00 $\pm$ 0.00 0.9586 $\pm$ 0.0761 0.0004 $\pm$ 0.0001 0.0109 $\pm$ 0.0046 0.0330 $\pm$ 0.0045
607 GPT-4o 1.00 $\pm$ 0.00 0.9931 $\pm$ 0.1001 0.0113 $\pm$ 0.0031 1.3643 $\pm$ 0.4998 0.0083 $\pm$ 0.0012
Gemini 3 Flash 1.00 $\pm$ 0.00 0.9754 $\pm$ 0.0724 0.0091 $\pm$ 0.0012 2.8852 $\pm$ 0.5663 0.0032 $\pm$ 0.0003
GPT-5.2 1.00 $\pm$ 0.00 0.9578 $\pm$ 0.0667 0.0090 $\pm$ 0.0012 1.5754 $\pm$ 0.3005 0.0057 $\pm$ 0.0003
Gemini 2.5 Flash 1.00 $\pm$ 0.00 0.9612 $\pm$ 0.0611 0.0085 $\pm$ 0.0011 2.4684 $\pm$ 0.5187 0.0034 $\pm$ 0.0004
Claude Opus 4.5 1.00 $\pm$ 0.00 1.0013 $\pm$ 0.1086 0.0056 $\pm$ 0.0011 0.1576 $\pm$ 0.0482 0.0358 $\pm$ 0.0016
GPT-5 1.00 $\pm$ 0.00 1.0077 $\pm$ 0.1079 0.0053 $\pm$ 0.0010 0.7238 $\pm$ 0.1735 0.0073 $\pm$ 0.0005
Kimi K2.5 1.00 $\pm$ 0.00 0.9899 $\pm$ 0.1223 0.0049 $\pm$ 0.0008 1.2700 $\pm$ 0.2678 0.0038 $\pm$ 0.0003
Claude Opus 4.6 1.00 $\pm$ 0.00 0.9798 $\pm$ 0.0845 0.0046 $\pm$ 0.0005 0.1372 $\pm$ 0.0153 0.0338 $\pm$ 0.0001
DeepSeek Chat 1.00 $\pm$ 0.00 0.9787 $\pm$ 0.0989 0.0036 $\pm$ 0.0003 2.0750 $\pm$ 0.2060 0.0017 $\pm$ 0.0000
GLM-5 1.00 $\pm$ 0.00 0.9850 $\pm$ 0.0631 0.0015 $\pm$ 0.0003 0.1741 $\pm$ 0.0428 0.0085 $\pm$ 0.0007
Gemini 3 Pro 1.00 $\pm$ 0.00 1.0523 $\pm$ 0.1248 0.0002 $\pm$ 0.0001 0.0060 $\pm$ 0.0029 0.0349 $\pm$ 0.0042
785 Gemini 3 Flash 0.98 $\pm$ 0.14 0.9794 $\pm$ 0.2044 0.0072 $\pm$ 0.0016 1.9708 $\pm$ 0.5443 0.0037 $\pm$ 0.0005
GPT-4o 0.90 $\pm$ 0.30 0.8845 $\pm$ 0.3341 0.0068 $\pm$ 0.0051 0.6551 $\pm$ 1.0487 0.0116 $\pm$ 0.0034
GPT-5.2 1.00 $\pm$ 0.00 0.5852 $\pm$ 0.0956 0.0067 $\pm$ 0.0012 0.8905 $\pm$ 0.2136 0.0075 $\pm$ 0.0008
Gemini 2.5 Flash 1.00 $\pm$ 0.00 0.9617 $\pm$ 0.1679 0.0066 $\pm$ 0.0012 1.5462 $\pm$ 0.3506 0.0043 $\pm$ 0.0003
Gemini 3 Pro 1.00 $\pm$ 0.00 1.0416 $\pm$ 0.1996 0.0050 $\pm$ 0.0040 0.1907 $\pm$ 0.2120 0.0265 $\pm$ 0.0084
Claude Opus 4.5 1.00 $\pm$ 0.00 0.9943 $\pm$ 0.1450 0.0046 $\pm$ 0.0006 0.1168 $\pm$ 0.0156 0.0392 $\pm$ 0.0002
Kimi K2.5 1.00 $\pm$ 0.00 0.9698 $\pm$ 0.1714 0.0040 $\pm$ 0.0009 0.8368 $\pm$ 0.2027 0.0048 $\pm$ 0.0004
Claude Opus 4.6 1.00 $\pm$ 0.00 0.9938 $\pm$ 0.1542 0.0038 $\pm$ 0.0007 0.1054 $\pm$ 0.0199 0.0365 $\pm$ 0.0001
GPT-5 0.96 $\pm$ 0.20 0.5326 $\pm$ 0.1370 0.0038 $\pm$ 0.0011 0.4073 $\pm$ 0.1385 0.0097 $\pm$ 0.0009
GLM-5 1.00 $\pm$ 0.00 0.7441 $\pm$ 0.2308 0.0011 $\pm$ 0.0003 0.1135 $\pm$ 0.0351 0.0101 $\pm$ 0.0009
DeepSeek Chat 0.26 $\pm$ 0.44 0.1633 $\pm$ 0.2789 0.0006 $\pm$ 0.0011 0.3938 $\pm$ 0.4810 0.0063 $\pm$ 0.0000
813 GPT-4o 0.98 $\pm$ 0.14 1.4618 $\pm$ 0.2675 0.0163 $\pm$ 0.0049 1.5711 $\pm$ 0.6311 0.0106 $\pm$ 0.0014
Gemini 3 Pro 1.00 $\pm$ 0.00 1.4871 $\pm$ 0.2484 0.0150 $\pm$ 0.0043 0.6948 $\pm$ 0.2329 0.0215 $\pm$ 0.0015
Gemini 3 Flash 1.00 $\pm$ 0.00 1.4672 $\pm$ 0.1450 0.0147 $\pm$ 0.0045 3.2074 $\pm$ 1.3323 0.0046 $\pm$ 0.0008
GPT-5.2 1.00 $\pm$ 0.00 1.4145 $\pm$ 0.1573 0.0147 $\pm$ 0.0019 1.8987 $\pm$ 0.3136 0.0077 $\pm$ 0.0010
Gemini 2.5 Flash 0.96 $\pm$ 0.20 1.3410 $\pm$ 0.3267 0.0119 $\pm$ 0.0031 2.2072 $\pm$ 0.7463 0.0056 $\pm$ 0.0007
Claude Opus 4.5 1.00 $\pm$ 0.00 1.5073 $\pm$ 0.1882 0.0111 $\pm$ 0.0018 0.2780 $\pm$ 0.0465 0.0398 $\pm$ 0.0001
Claude Opus 4.6 1.00 $\pm$ 0.00 1.4065 $\pm$ 0.1678 0.0086 $\pm$ 0.0010 0.2246 $\pm$ 0.0259 0.0381 $\pm$ 0.0001
Kimi K2.5 0.94 $\pm$ 0.24 1.3564 $\pm$ 0.4013 0.0083 $\pm$ 0.0026 1.5839 $\pm$ 0.7307 0.0055 $\pm$ 0.0013
GPT-5 1.00 $\pm$ 0.00 1.4285 $\pm$ 0.1680 0.0081 $\pm$ 0.0011 0.7875 $\pm$ 0.1343 0.0102 $\pm$ 0.0007
GLM-5 1.00 $\pm$ 0.00 1.5179 $\pm$ 0.1824 0.0017 $\pm$ 0.0009 0.1106 $\pm$ 0.0876 0.0153 $\pm$ 0.0039
DeepSeek Chat 0.18 $\pm$ 0.38 0.2602 $\pm$ 0.5612 0.0011 $\pm$ 0.0024 0.8861 $\pm$ 1.2098 0.0071 $\pm$ 0.0000
895 GPT-4o 1.00 $\pm$ 0.00 0.9521 $\pm$ 0.0391 0.0731 $\pm$ 0.0134 5.6117 $\pm$ 1.6328 0.0130 $\pm$ 0.0034
Gemini 2.5 Flash 1.00 $\pm$ 0.00 0.9775 $\pm$ 0.0408 0.0582 $\pm$ 0.0089 10.1518 $\pm$ 2.6428 0.0057 $\pm$ 0.0009
Gemini 3 Flash 1.00 $\pm$ 0.00 0.9624 $\pm$ 0.0633 0.0542 $\pm$ 0.0071 10.7101 $\pm$ 2.2685 0.0051 $\pm$ 0.0005
GPT-5 0.94 $\pm$ 0.24 0.9035 $\pm$ 0.2332 0.0374 $\pm$ 0.0126 2.7033 $\pm$ 1.1059 0.0147 $\pm$ 0.0017
Gemini 3 Pro 1.00 $\pm$ 0.00 0.9835 $\pm$ 0.0725 0.0325 $\pm$ 0.0062 1.2292 $\pm$ 0.5011 0.0265 $\pm$ 0.0071
Claude Opus 4.5 1.00 $\pm$ 0.00 0.9504 $\pm$ 0.0780 0.0325 $\pm$ 0.0029 0.6086 $\pm$ 0.0552 0.0534 $\pm$ 0.0005
Claude Opus 4.6 1.00 $\pm$ 0.00 0.9633 $\pm$ 0.0758 0.0314 $\pm$ 0.0021 0.6513 $\pm$ 0.0443 0.0482 $\pm$ 0.0003
Kimi K2.5 1.00 $\pm$ 0.00 0.9716 $\pm$ 0.0766 0.0239 $\pm$ 0.0041 3.8241 $\pm$ 1.1679 0.0063 $\pm$ 0.0014
GLM-5 1.00 $\pm$ 0.00 0.9824 $\pm$ 0.0487 0.0073 $\pm$ 0.0018 0.5158 $\pm$ 0.1935 0.0141 $\pm$ 0.0024
DeepSeek Chat 0.28 $\pm$ 0.45 0.2648 $\pm$ 0.4250 0.0059 $\pm$ 0.0098 3.2359 $\pm$ 3.9557 0.0065 $\pm$ 0.0007
GPT-5.2 0.12 $\pm$ 0.32 0.1108 $\pm$ 0.3007 0.0049 $\pm$ 0.0132 0.2930 $\pm$ 0.8119 0.1386 $\pm$ 0.0007
968 GPT-4o 1.00 $\pm$ 0.00 0.9189 $\pm$ 0.1193 0.0034 $\pm$ 0.0009 0.3918 $\pm$ 0.1327 0.0087 $\pm$ 0.0010
Claude Opus 4.5 1.00 $\pm$ 0.00 0.9280 $\pm$ 0.1291 0.0030 $\pm$ 0.0005 0.1208 $\pm$ 0.0227 0.0250 $\pm$ 0.0006
Gemini 2.5 Flash 1.00 $\pm$ 0.00 0.9239 $\pm$ 0.1139 0.0028 $\pm$ 0.0004 0.8190 $\pm$ 0.1684 0.0035 $\pm$ 0.0004
Gemini 3 Flash 1.00 $\pm$ 0.00 0.8686 $\pm$ 0.1100 0.0028 $\pm$ 0.0004 0.8079 $\pm$ 0.1836 0.0034 $\pm$ 0.0005
Claude Opus 4.6 1.00 $\pm$ 0.00 0.9534 $\pm$ 0.1726 0.0022 $\pm$ 0.0006 0.0727 $\pm$ 0.0182 0.0308 $\pm$ 0.0060
Gemini 3 Pro 1.00 $\pm$ 0.00 0.9264 $\pm$ 0.1033 0.0019 $\pm$ 0.0002 0.1159 $\pm$ 0.0205 0.0167 $\pm$ 0.0016
Kimi K2.5 0.96 $\pm$ 0.20 0.8471 $\pm$ 0.2204 0.0019 $\pm$ 0.0008 0.5379 $\pm$ 0.5124 0.0037 $\pm$ 0.0004
GPT-5 0.96 $\pm$ 0.20 0.8096 $\pm$ 0.2532 0.0017 $\pm$ 0.0006 0.2090 $\pm$ 0.0933 0.0083 $\pm$ 0.0010
DeepSeek Chat 0.74 $\pm$ 0.44 0.6497 $\pm$ 0.3990 0.0008 $\pm$ 0.0005 0.4673 $\pm$ 0.2764 0.0023 $\pm$ 0.0000
GLM-5 0.98 $\pm$ 0.14 0.9016 $\pm$ 0.1775 0.0003 $\pm$ 0.0001 0.0236 $\pm$ 0.0119 0.0114 $\pm$ 0.0021
GPT-5.2 0.04 $\pm$ 0.20 0.0352 $\pm$ 0.1723 0.0001 $\pm$ 0.0006 0.0212 $\pm$ 0.1134 0.1509 $\pm$ 0.0000
Appendix E TPC-H evaluation
E.1. Model and Query Selection
We selected three last generation frontier models available at the time of our original experiments in January 2026: Gemini 3 Pro, Claude Opus 4.5, and GPT-5.2. We excluded additional models due to budget constraints, as a single TPC-H query run at the specified scale factor costs approximately \$1 within our proposed EMR cluster. Given 50 replicas per model and the orginal set of 10 models from the BIRD evaluation, the total cost for four TPC-H queries would reach roughly \$2,000, excluding lower scale factor tests. Because an exhaustive assessment of all available models and queries would not further clarify the primary contributions of this paper, we limited this evaluation to a representative subset.
Regarding query selection, we prioritized TPC-H queries that Claude Opus 4.5 could not translate correctly. We then verified that (1) accuracy remained imperfect in the other two models and (2) no model had memorized the TPC-H specification during training. While methods exist to prove memorization in production Large Language Models (LLMs), verifying the absolute absence of memorization remains an open challenge. As a workaround, we adapted the methodology from (ahmed2026extractingbooksproductionlanguage), implementing the code at https://github.com/GEizaguirre/memorization-LLM-prod, and attempted to detect TPC-H memorization in the models; however, these efforts were unsuccessful.
Appendix F Detailed Results
Our TPC-H evaluation averages each metric over 50 iterations at scale factor 1. Table LABEL:tab:appendix-tpch-sf1-queries provides a detailed breakdown of the resulting Text-to-SQL and Text-to-Big SQL metrics for each query and model.
Table 5. Per-query TPC-H (SF1) metrics by model (mean $\pm$ std across runs). VES, VES*, and VCES follow the paper definitions; SVQ reports the expected cost per valid query.
Query ID Model EX VES VES* VCES SVQ
1 GPT-5.2 0.60 $\pm$ 0.49 0.6507 $\pm$ 0.5318 0.2608 $\pm$ 0.2134 23.1794 $\pm$ 19.0188 0.0188 $\pm$ 0.0000
Claude Opus 4.5 1.00 $\pm$ 0.00 0.9567 $\pm$ 0.1057 0.2513 $\pm$ 0.0138 3.6030 $\pm$ 0.2671 0.0697 $\pm$ 0.0015
Gemini 3 Pro (T) 0.00 $\pm$ 0.00 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
17 Claude Opus 4.5 0.20 $\pm$ 0.40 0.2028 $\pm$ 0.4235 0.0323 $\pm$ 0.0647 0.3865 $\pm$ 0.6072 0.4182 $\pm$ 0.0025
Gemini 3 Pro (T) 0.10 $\pm$ 0.30 0.1294 $\pm$ 0.3883 0.0140 $\pm$ 0.0420 0.3525 $\pm$ 0.7373 0.3973 $\pm$ 0.0000
GPT-5.2 0.40 $\pm$ 0.49 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0168 $\pm$ 0.0004
18 Gemini 3 Pro (T) 1.00 $\pm$ 0.00 1.3223 $\pm$ 0.1182 0.2449 $\pm$ 0.0232 6.4131 $\pm$ 0.8326 0.0382 $\pm$ 0.0019
GPT-5.2 0.40 $\pm$ 0.49 0.2206 $\pm$ 0.6619 0.0711 $\pm$ 0.2133 7.5480 $\pm$ 17.8560 0.0236 $\pm$ 0.0030
Claude Opus 4.5 0.00 $\pm$ 0.00 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
21 Claude Opus 4.5 0.80 $\pm$ 0.40 0.7732 $\pm$ 0.3889 0.2639 $\pm$ 0.1321 2.8673 $\pm$ 1.7429 0.1150 $\pm$ 0.0003
GPT-5.2 0.20 $\pm$ 0.40 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0915 $\pm$ 0.0000
Gemini 3 Pro (T) 0.00 $\pm$ 0.00 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
We then execute the generated SQL queries across different scale factors and average the proposed metrics for all queries. These results are shown in Table LABEL:tab:appendix-tpch-scale-factors.
Table 6. TPC-H metrics by query, model, and scale factor (mean $\pm$ std across runs).
Query ID Scale Factor Model VES VES* VCES SVQ
1 10 Gemini 3 Pro 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
Opus 4.5 0.5668 $\pm$ 0.0000 0.2133 $\pm$ 0.0064 1.8351 $\pm$ 0.0783 0.1163 $\pm$ 0.0015
GPT-5.2 0.3235 $\pm$ 0.2641 0.1857 $\pm$ 0.1517 3.0724 $\pm$ 2.5113 0.1007 $\pm$ 0.0000
100 Gemini 3 Pro 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
Opus 4.5 0.2654 $\pm$ 0.0000 0.1494 $\pm$ 0.0031 0.8809 $\pm$ 0.0259 0.1697 $\pm$ 0.0015
GPT-5.2 0.1599 $\pm$ 0.1306 0.1170 $\pm$ 0.0955 1.0526 $\pm$ 0.8598 0.1852 $\pm$ 0.0000
1000 Gemini 3 Pro 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
Opus 4.5 0.0484 $\pm$ 0.0000 0.0424 $\pm$ 0.0002 0.0683 $\pm$ 0.0006 0.6199 $\pm$ 0.0015
GPT-5.2 0.0289 $\pm$ 0.0236 0.0271 $\pm$ 0.0221 0.0478 $\pm$ 0.0390 0.9439 $\pm$ 0.0000
17 10 Gemini 3 Pro 0.0519 $\pm$ 0.1556 0.0121 $\pm$ 0.0362 0.1148 $\pm$ 0.3444 0.8777 $\pm$ 0.0000
Opus 4.5 0.1040 $\pm$ 0.2079 0.0284 $\pm$ 0.0570 0.1814 $\pm$ 0.3637 0.6704 $\pm$ 0.0026
GPT-5.2 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.1025 $\pm$ 0.0004
100 Gemini 3 Pro 0.0332 $\pm$ 0.0997 0.0107 $\pm$ 0.0320 0.0807 $\pm$ 0.2422 1.1488 $\pm$ 0.0000
Opus 4.5 0.0669 $\pm$ 0.1337 0.0247 $\pm$ 0.0494 0.1335 $\pm$ 0.2674 0.8111 $\pm$ 0.0026
GPT-5.2 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.1357 $\pm$ 0.0004
1000 Gemini 3 Pro 0.0051 $\pm$ 0.0154 0.0039 $\pm$ 0.0116 0.0071 $\pm$ 0.0213 5.2761 $\pm$ 0.0000
Opus 4.5 0.0094 $\pm$ 0.0187 0.0075 $\pm$ 0.0151 0.0113 $\pm$ 0.0225 3.2374 $\pm$ 0.0026
GPT-5.2 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.7089 $\pm$ 0.0004
18 10 Gemini 3 Pro 0.5006 $\pm$ 0.0000 0.1879 $\pm$ 0.0142 1.8557 $\pm$ 0.1625 0.1014 $\pm$ 0.0019
Opus 4.5 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
GPT-5.2 0.1017 $\pm$ 0.3052 0.0517 $\pm$ 0.1551 1.0504 $\pm$ 3.1511 0.1172 $\pm$ 0.0029
100 Gemini 3 Pro 0.2563 $\pm$ 0.0000 0.1382 $\pm$ 0.0080 0.8541 $\pm$ 0.0561 0.1619 $\pm$ 0.0019
Opus 4.5 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
GPT-5.2 0.0616 $\pm$ 0.1849 0.0389 $\pm$ 0.1166 0.5274 $\pm$ 1.5823 0.1783 $\pm$ 0.0029
1000 Gemini 3 Pro 0.0481 $\pm$ 0.0000 0.0414 $\pm$ 0.0008 0.0592 $\pm$ 0.0012 0.6997 $\pm$ 0.0019
Opus 4.5 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
GPT-5.2 0.0187 $\pm$ 0.0562 0.0159 $\pm$ 0.0477 0.0737 $\pm$ 0.2211 0.5334 $\pm$ 0.0029
21 10 Gemini 3 Pro 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
Opus 4.5 0.3922 $\pm$ 0.1961 0.1984 $\pm$ 0.0992 1.2211 $\pm$ 0.6107 0.2233 $\pm$ 0.0003
GPT-5.2 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.4713 $\pm$ 0.0000
100 Gemini 3 Pro 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
Opus 4.5 0.2300 $\pm$ 0.1150 0.1462 $\pm$ 0.0731 0.6527 $\pm$ 0.3264 0.3002 $\pm$ 0.0003
GPT-5.2 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.7582 $\pm$ 0.0000
1000 Gemini 3 Pro 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000
Opus 4.5 0.0308 $\pm$ 0.0154 0.0286 $\pm$ 0.0143 0.0241 $\pm$ 0.0121 1.5039 $\pm$ 0.0003
GPT-5.2 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 4.4840 $\pm$ 0.0000